# Brain-inspired Large Language Model

### v2

@@ 4 Experiments
We released BriLLM-Chinese-V0.5 model.
@@
We use the subset from the Chinese version of Wikipedia, which contains over 100M Chinese characters. We truncate the long sentences into small sentences with a maximum length of 16. We select a vocabulary of 4,000 tokens consisting of the most frequently used Chinese characters.
@@
BriLLM is implemented using PyTorch. It uses sine and cosine positional encoding, GeLU as the activation function, cross-entropy loss for next-token prediction, and an embedding size of $d_{model}=32$. We used the AdamW optimizer with $\beta_{1}=0.9$, $\beta_{2}=0.999$ and $\epsilon=10^{-8}$. The model size is about $512+4000*4000*(32*32+32)\approx 16B$. We trained our models on one machine with 8 NVIDIA A800 GPUs for 1.5k steps. The training loss is shown in Figure 4.

### v3

@@ 4 Experiments
We released BriLLM-Chinese and BriLLM-English model.
@@
For BriLLM-Chinese and BriLLM-English, we use the Chinese and English versions of Wikipedia respectively, each containing over 100M tokens. We truncate the long sentences into small sentences with a maximum length of 32. We select a vocabulary of 4,000 tokens consisting of the most frequently used Chinese characters.
Sparse Training
BriLLM is a sparse model, where the occurrence probability of most bigrams is very low or even zero, allowing us to leverage this characteristic for sparse training. We set the connection weights corresponding to low-frequency bigrams (those not appearing in the training set) to be shared and update them randomly. This approach reduces the model size to approximately 10% of the original while significantly accelerating the training speed.
@@
BriLLM is implemented using PyTorch. It uses sine and cosine positional encoding, GeLU as the activation function, cross-entropy loss for next-token prediction, and an embedding size of $d_{model}=32$. We used the AdamW optimizer with $\beta_{1}=0.9$, $\beta_{2}=0.999$ and $\epsilon=10^{-8}$. The original model size is about $512+4000*4000*(32*32+32)\approx 16B$, but after applying sparse training, the actual size is reduced to around $2B$. We trained our models on one machine with 8 NVIDIA A800 GPUs for 1.5k steps. The training loss is shown in Figure 4.